Please help me solve this equation:
Solve the system of equations.
X - Y + 2z = -3
X + 2y + 3z = 4
2x + y + z = -3


Thank you for your time I really appreciated

Answer by Alan3354(69443)

Please help me solve this equation:
Solve the system of equations.
X - Y + 2z = -3
X + 2y + 3z = 4
2x + y + z = -3
----------------
Subtract eqns 1 & 2, that eliminates x.
Multiply eqn 1 by 2, then subtract eqn 3.
That eliminates x, leaving 2 eqns in y & z.
Can you do the rest?
email via the TY note for help or to check your work.